OpenStudy (anonymous):

An astronaut weighs 170 pounds on Earth but only 68 pounds on Mars. While visiting the planet, the astronaut removes a rock sample that weighs 25 pounds on Mars. What is the weight of the rock sample on Earth? Help me solve.

OpenStudy (anonymous):

@amistre64 @heather040200 @marissalovescats @loujoelou

OpenStudy (chmvijay):

68 pounds on mars = 170 pound on earth then 25 pounds on earth =170*25 /68 =62.5

OpenStudy (anonymous):

but that's not one of my choices.

OpenStudy (chmvijay):

what is ur choice

OpenStudy (anonymous):

a. 0.22 b. 4 c. 0.5 d. not enough info.

OpenStudy (anonymous):

I need to understand HOW to do it because if I don't I'd be lost on ALL my other questions.. & It's really important I pass this..

OpenStudy (chmvijay):

the units are in pounds only right

OpenStudy (amistre64):

mg1 = 170 m = 170/g1 , what is gravity in terms of lbs? i forget the ratio mg2 = 68 m = 68/g2 170/g1 = 68/g2 ; therefore g2 = 68g1/170

OpenStudy (anonymous):

yes, I am guessing because it says pounds.

OpenStudy (anonymous):

@amistre64 what?

OpenStudy (amistre64):

theres enough info if you can recall g1 in proper terms

OpenStudy (anonymous):

okay, can one of you please just explain to me how to do this? :) please.

OpenStudy (amistre64):

F = ma, where a is the acceleration of gravity on the planet this is usually expressed in terms of Newtons; or kg-m/sec^2

OpenStudy (anonymous):

@chmvijay Thank you, I didn't notice you explained thanks.

OpenStudy (amistre64):

1 newton is equal to .2248 lbs per foot

OpenStudy (anonymous):

That was a choice sorry whoops. :)

OpenStudy (chmvijay):

LOL what was choice

OpenStudy (anonymous):

LOL, yeah thanks !

OpenStudy (chmvijay):

but i didnt get what ur options are wrong u inserted here

OpenStudy (anonymous):

I know because I was looking at the wrong answer choices.

OpenStudy (anonymous):

An astronaut weighs 170 pounds on Earth and 400 pounds on Jupiter. The astronaut captures an alien that weighs 120 pounds on Jupiter. What is the alien's weight on Earth?

OpenStudy (chmvijay):

hey @amistre64 even if u calculate by that method also the same answer you will get :) i got ur point u wanted to find gravitation on mars bases on other :)

OpenStudy (anonymous):

would that be 283?

OpenStudy (chmvijay):

nope

OpenStudy (anonymous):

400 lbs = 170 lbs 120 lbs = 400x120/170 = 282.3529.... rounded 283.@chmvijay

OpenStudy (anonymous):

help please. :(

OpenStudy (chmvijay):

400 weighs on Jupiter =170 on earth 120 weight on Jupiter =170*120 /400

OpenStudy (anonymous):

Thats the way to solve these questions? That method always?

OpenStudy (chmvijay):

yup :)

OpenStudy (chmvijay):